OpenStudy (anonymous):

rationalize the denominator √3/(√17) +3

OpenStudy (anonymous):

Is it this? \(\large \frac{\sqrt{3}}{\sqrt{17} + 3}\)

OpenStudy (anonymous):

yeah the 3 isnt under the square root.

OpenStudy (angela210793):

multiply with sqr17-3 up and down sqr3(sqr17-3)/8=sqr51-3sqr3

OpenStudy (anonymous):

oh, then no.

OpenStudy (angela210793):

it isn't? :/

OpenStudy (anonymous):

Just put the two numbers over a common denominator, then add them, then rationalize the denominator.

OpenStudy (anonymous):

what two numbers?

OpenStudy (anonymous):

\(\frac{\sqrt{3}}{\sqrt{17}}\) and 3. The two numbers you're adding. They need to have a common denominator in order to add them.

OpenStudy (anonymous):

there is no common denominator.

OpenStudy (anonymous):

Sure there is. The LCD of 1 and sqrt(17) is sqrt(17)

OpenStudy (angela210793):

thr is

OpenStudy (anonymous):

im confused...

OpenStudy (anonymous):

\[3 = \frac{3}{1} = \frac{3}{1} \cdot 1 = \frac{3}{1} \cdot \frac{\sqrt{17}}{\sqrt{17}} = \frac{3\sqrt{17}}{\sqrt{17}}\]

OpenStudy (anonymous):

So then you can add it

OpenStudy (anonymous):

so it would be 3√20 / √17 ?

OpenStudy (anonymous):

No, you cannot add sqrts.

OpenStudy (anonymous):

\[\sqrt{3} + 3\sqrt{17}\] cannot be futher combined.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i am wiling to bet that this is the question\[\large \frac{\sqrt{3}}{\sqrt{17} + 3}\]

OpenStudy (anonymous):

I thought so, but he said no.

OpenStudy (anonymous):

because of the instructions that that "rationalize the denominator"

OpenStudy (anonymous):

actually i think he said yeah the 3 isnt under the square root.

OpenStudy (anonymous):

oh, crud.

OpenStudy (anonymous):

yeah, the three isnt under the square root.

OpenStudy (anonymous):

\[\large \frac{3}{\sqrt{17} + 3}\]

OpenStudy (anonymous):

I thought he said it wasn't under the fraction.

OpenStudy (anonymous):

In that case, ignore what I said. You just want to multiply top and bottom by the conjugate of sqrt(17) + 3

OpenStudy (anonymous):

@xxcrash do you see the correct question anywhere here?

OpenStudy (anonymous):

no i dont understand how to do it

OpenStudy (anonymous):

wouldnt the answer be √51 over 20? because 3(17) is 51? and then leave 3 on the bottom and add it to 17?

OpenStudy (anonymous):

\[\frac{3}{\sqrt{17}+3}\times \frac{\sqrt{17}-3}{\sqrt{17}-3}\] \[\frac{3(\sqrt{17}-3)}{\sqrt{17}\times \sqrt{17}-9}\] \[\frac{3(\sqrt{17}-3)}{17-9}\] \[\frac{3(\sqrt{17}-3)}{8}\]

OpenStudy (anonymous):

the 3 on top is under a √ too...

OpenStudy (anonymous):

then just rewrite exactly what i wrote, replacing the 3 by a root 3

OpenStudy (anonymous):

but then wouldnt u multiply √3 by √17 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

why ?

OpenStudy (anonymous):

there is nothing in that expression that indicates for you to multiply \[\sqrt{3}\times \sqrt{17}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lets go slow and be clear. first of all, is this what you started with \[ \frac{\sqrt{3}}{\sqrt{17} + 3}\]

OpenStudy (anonymous):

yeah